Using the same economic data, you have been asked to recommend the divestment of one of the SBUs and purchase of Company D with expected returns of Boom 18%, Normal 10%, Decline 2%. The % of portfolio for the new investment would be the same as the one sold. Which company do you suggest selling? What will be the portfolio return now?

Scenario   

Probability

Return A

Return B

Return C

Boom

0.2

0.2

0.3

0.45

Normal

0.6

0.12

0.13

0.12

Decline

0.2

0

-0.1

-0.40

% of portfolio

0.4

0.35

0.25
Answer #1

Expected return=Sum(probability*return)

A:
=0.2*0.2+0.6*0.12+0.2*0=0.112

B:
=0.2*0.3+0.6*0.13+0.2*(-0.1)=0.118

C:
=0.2*0.45+0.6*0.12+0.2*(-0.4)=0.082

D:
=0.2*0.18+0.6*0.1+0.2*0.02=0.1

Sell C and buy D

Portfolio return=0.4*0.112+0.35*0.118+0.25*0.1=11.11%